During a baseball game, a batter hits a pop-up. The ball goes straight upwards into the air and then comes back down to the ground in 6.0 seconds, calculate how fast the ball came off the bat and how high the ball went.
Answer #1

time it take to reach max height

t = 6/2 = 3 s

velocity with which it comes or with which it was projected upwards

v = gt = 9.8* 3 = 29.4 m/s

=======

max height it reached

h = v^2 / 2 g

h = 29.4^2 / 19.6

h = 44.1 m

=========

comment in case any doubt, will reply for sure.. goodluck